Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2017-09-13

Observers: Mariusz Gromadzki, Kris Rybicki
TAT: Joe Anderson, Annalisa De Cia
Support team: Morgan Fraser, Emma Callis

Calibrations: biases, dome arcs+flats,

Conditions: Clouds and high humidity at the first hour of the night, then humidity decrease to 23%  seeing around 1.3 


UT       Target            Type            Comment
==============================================================================================
00:12    ATLAS2017jrp        class         Gr13 silt=1.5
00:34    AT2017grw           class         Gr13 galaxy - maybe wrong object?
00:55    EG274               standard 
01:33    AT2017fwm           followup      Gr13          slit=1.0
02:28    SMT17scn            followup      Gr11&Gr16
04:25    AT2017gpp           followup      Gr11&Gr16   
06:27    AT2017gsa           class         Gr13         CV in outburst? 
06:50    Feige110            standard      exposure without guiding because of guider problem
         Technical problem with guider     10min lost
07:10    EFOSC -> SOFI
         Technical problem with guider & SOFI detector   1h 10min (50min+20min) lost
08:30    Feige110            standard     
08:40    HIP021545           telluric     TEL_AT2017gax
08:53    AT2017gax           followup     BG + JHK
09:35    HIP029337           telluric     TEL_AT2017fzw
09:57    AT2017fzw           followup     BG        




Comments